OpenStudy (dls):

Help with 3rd limit problem~

OpenStudy (dls):

\[\lim_{x \rightarrow \infty } \sqrt{x}(\sqrt{x+c}-\sqrt{x})\]

OpenStudy (dls):

I'm here